# Detecting Cultural Differences in News Video Thumbnails via Computational Aesthetics

### v1

@@ 1 Introduction
Visual imagery is an extremely important channel for international news coverage, as it connects the audience to events that may otherwise be removed from direct experience, and serves as a powerful tool for conveying emotion (Smith and Joffe 2013). For threatening events, such as global warming, wars, or pandemics, images have also been shown to be significant drivers of public engagement, “providing viewers with tangible and emotion-evoking examples that act as visual proof” (Joffe 2008; Martikainen and Sakki 2021). As such, highly visual mediums like YouTube play a significant role in modern news consumption (Burgess and Green 2018; Xie et al. 2011).
@@
Since statistical image properties are highly dependent on the content of the image (Kao, He, and Huang 2017; Machajdik and Hanbury 2010), our proposed framework first clusters images about a given news story by “visual themes” (i.e. different aspects or sub-events), We then analyze the aesthetic differences between channels. Lastly, we consider the views, likes, and comments for each video, in order to capture potential cross-cultural differences in how thumbnail properties may influence video performance.
@@ 2.1 Computational Aesthetics
In automatic image aesthetic assessment, also known as computational aesthetics, a model is trained to quantify the beauty of an image through some aesthetic score. Often, these models assess an image by measuring the degree to which it adheres to fundamental but abstract principles of photography, such as balance, rhythm, harmony, unity, tone, etc. (Anwar et al. 2021).
@@
These rely on extracting pre-defined low-level features from images and fitting a machine learning model, such as SVM, to do the prediction. Examples of features that have been extensively explored previously include sharpness, colorfulness, contrast, and texture (Lo, Liu, and Chen 2013; Redi et al. 2015; Aydın, Smolic, and Gross 2014).
@@
These leverage neural networks to learn low and high-level image features without (or with relatively little) human instruction. These are trained on large image datasets such as Photo.net (Datta et al. 2006), AVA (Murray, Marchesotti, and Perronnin 2012), and CUHK-PQ (Luo, Wang, and Tang 2011), and the specific architecture can vary. Doshi et al. (Doshi, Shikkenawis, and Mitra 2020) fine-tuned two pre-trained models (AlexNet and VGG) to classify images into high and low aesthetic categories. Kao et al. (Kao, He, and Huang 2017) investigated using a Multi-Task Convolutional Neutral network (MTCNN) to jointly perform aesthetic assessment and semantic recognition.
@@
In addition to predicting image aesthetics such as quality, beauty, and liking (Bartho, Thoemmes, and Redies 2023; Peng and JEMMOTT III 2018), there is evidence to suggest that computational aesthetics can serve as a valuable starting point for a variety of downstream tasks.
For instance, one experiment demonstrated the ability to infer the personality traits (“Big Five Inventory”) of Flickr users based on the color, composition, texture, and face features of images that the user tagged as ‘favorites’ on the platform (Segalin et al. 2016).
Likewise, Machajdik et al. (Machajdik and Hanbury 2010) achieved state-of-the-art results on image emotion classification using a similar set of features, including color, texture, composition, and content, inspired by psychology and art theory. However, results were not uniform: content features such as the number of faces and relative size of the biggest face were good discriminators between fear/disgust vs. amusement in the International Affective Picture System dataset, whereas color features was more predictive in an artistic photography dataset collected from the DeviantArt website.
@@ 2.2 Image Analysis in News Media
One investigation analyzed images of the COVID-19 pandemic appearing in Finnish newspapers, focusing on the social representations of individuals from varying age groups (Martikainen and Sakki 2021). The authors noted that the angle of the photo often contributed to visual rhetoric, for instance, images of children studying at home emphasized a high angle, placing the spectator in the authoritative position overseeing the child whereas those of young adults tended to be taken at eye-level.
Similarly, in a study on how the media covered Pope Francis’ visit to Cuba, Thomson et al. discovered that photo angles in U.S. sources framed the Pope at higher positions than Cuban politicians whereas local sources used a more equal leveling (Thomson, Perreault, and Duffy 2018). Another contrast was the content of the images, with Cuban media focusing almost exclusively on the Pope, whereas U.S. media more often published photographs of Cuban people.  (Fahmy 2010; Rafiee, Spooren, and Sanders 2023).
@@
Although these frequently involve analyzing text data to compare sentiments and topics (Wallbing 2021; Chen et al. 2023), the visual medium of news media has also been explored. Bhargava et al. (Bhargava, Bishop, and Zuckerman 2020) proposed Photo Spaces, a visualization tool to describe visual narratives for an event, based on image embeddings produced by a pre-trained ResNet-50 model. The spatial positioning of images in the photo-space is determined by reducing these embeddings into two dimensions, via the Uniform Manifold Approximation and Projection technique, from which clusters corresponding to visual themes can be extracted. The authors document an example of the tool (See Figure 1) on abortion rights as a topic, adding the political affinity of sources (right- vs. left-leaning) through the color of each image’s border.
Figure 1: An example of Photo-Space visualization (Bhargava, Bishop, and Zuckerman 2020), where physical placement captures thematic relations, and image borders are color-coded by political affinity.
Similarly, Zhang et al. (Zhang and Peng 2022) evaluated the effectiveness of bag-of-visual-words, self-supervised learning, and transfer learning to cluster unlabeled image data. They conclude that transfer learning using pre-trained models to obtain embeddings outperforms the other methods, though the choice of the pre-trained model significantly affected clustering results.
@@ 2.3 Video Thumbnails
Thumbnails are small image representations of full-length videos. Thumbnails “tell the users what the video is about (i.e., be informative), and grab potential viewers’ attention (i.e., be visually appealing)” (Koh and Cui 2022).
Kim et al. (Kim et al. 2023) evaluated the design of thumbnails for data-driven news articles, concluding similarly that for thumbnails involving informative graphics, users prefer ones that are interpretable and attractive. For branded content, a thumbnail’s success (measured in view-through) can also be attributed to visual features, such as colorfulness and brightness (Koh and Cui 2022).
@@ 3.1 Data Collection
Inspired by previous work (Martikainen and Sakki 2021), we selected the COVID-19 pandemic as one of the international news events to examine. The complexity of the event gives rise to a diverse set of event themes—medical, economic, political—which may transfer to detectable visual themes. For similar reasons, we also selected the war in Ukraine, a more recent event with different event themes—military conflict, the humanitarian crisis, and geopolitics.
@@ 3.2 Visual Theme Identification
After some experimentation, our best results were obtained by using Concept (Grootendorst 2021), a CLIP and BERTopic-based open source tool that performs topic modeling on images, by first encoding images using CLIP, and then identifying clusters on top of those embeddings. An example of the results produced using this method is shown in Figure 2.
@@
Thumbnail clusters themselves carry no easily interpretable meaning, so to understand the visual themes being extracted we associated each cluster with five textual tags. We leveraged Tag2Text (Huang et al. 2023), a model capable of recognizing 3,429 commonly human-used categories, to generate tags for every image. Then, on a cluster level, we aggregated these tags to generate five tags that are most representative of the cluster’s theme.
@@ 3.3 Image Feature Extraction
Image features are extracted using a hand-crafted approach, where various pre-defined statistical measures are computed for each image. An exhaustive reference list of the 21 features, in eight feature groups, is displayed in Table 2. The majority of our feature set is adopted from Bartho et al. (Bartho, Thoemmes, and Redies 2023). A brief justification of each feature is included here, but the reader is encouraged to consult the references for the computational details and tuning parameters of each algorithm.
@@
Color (Datta et al. 2006), (Li and Chen 2009), (Mallon, Redies, and Hayn-Leichsenring 2014), (Schifanella, Redi, and Aiello 2015), (Thieleking et al. 2020), (Iigaya et al. 2021), (Geller et al. 2022) 5 Hue mean of hue channel in HSV space
@@
Dimension (Datta et al. 2006), (Iigaya et al. 2021) (Mallon, Redies, and Hayn-Leichsenring 2014) 2 Aspect Ratio image width divided by image height (pixels)
@@
Lightness (Schifanella, Redi, and Aiello 2015), (Sidhu et al. 2018), (Mather 2018), (Iigaya et al. 2021), (Peli 1990) 3 Contrast standard deviation of L* channel in L*a*b* space
@@
HOG (Bosch, Zisserman, and Munoz 2007), (Braun et al. 2013) 3 Self-Similarity similarity of HOG features
@@
Fourier (Redies et al. 2008) 2 Fourier Slope
@@
Symmetry (Brachmann and Redies 2016) (Bertamini and Rampone 2019) 2 Symmetry-LR
@@
CNN (Brachmann, Barth, and Redies 2017) 2 Sparseness
@@
Other (Savardi et al. 2018), (Zhang and Peng 2022) >1 Shot Scale close, medium, or long shot based on fine-tuned CNN
@@
Hue, Saturation, Lab(a), Lab(b), Color Entropy. The use of colors can affect the perceived mood of an image. Although there are multiple ways to represent colors numerically, the HSV space (for hue, saturation, and value), and the L*a*b* space (for lightness, green-red, and blue-yellow) both have been shown to be more aligned with the human perception of color (Fadaei 2021). Color-entropy captures the visual percept of the “colorfulness” of an image.
@@ 4 Validations
Using the PlacesCNN model (Zhou et al. 2014), which can label each image as indoor/outdoor, we measured the entropy of each theme using the Gini index of statistical dispersion.
@@ 5 Results
[…]
@@ 7 Application to Disinformation
Previous investigation has shown that multimodal disinformation that uses a text-plus-visual format is perceived as more credible than pure text (Hameleers et al. 2020). Visual disinformation can involve visual content that is deliberately manipulated or generated, or real but used out of context, or real but used in tandem with false claims (Cao et al. 2020). This paper focuses on the first two.

### v2

@@ 1 Introduction
Visual imagery is an extremely important channel for international news coverage, as it connects the audience to events that may otherwise be removed from direct experience, and serves as a powerful tool for conveying emotion (smith2013public). For threatening events, such as global warming, wars, or pandemics, images have also been shown to be significant drivers of public engagement, “providing viewers with tangible and emotion-evoking examples that act as visual proof” (joffe2008power; martikainen2021newspaper). As such, highly visual mediums like YouTube play a significant role in modern news consumption (burgess2018youtube; xie2011visual).
@@
Since statistical image properties are highly dependent on the content of the image (kao2017deep; machajdik2010affective), our proposed framework first clusters images about a given news story by “visual themes” (i.e. different aspects or sub-events), We then analyze the aesthetic differences between channels. Lastly, we consider the views, likes, and comments for each video, in order to capture potential cross-cultural differences in how thumbnail properties may influence video performance.
@@ 2.1 Computational Aesthetics
In automatic image aesthetic assessment, also known as computational aesthetics, a model is trained to quantify the beauty of an image through some aesthetic score. Often, these models assess an image by measuring the degree to which it adheres to fundamental but abstract principles of photography, such as balance, rhythm, harmony, unity, tone, etc. (anwar2021survey).
@@
These rely on extracting pre-defined low-level features from images and fitting a machine learning model, such as SVM, to do the prediction. Examples of features that have been extensively explored previously include sharpness, colorfulness, contrast, and texture (lo2013intelligent; redi2015beauty; aydin2014automated).
@@
These leverage neural networks to learn low and high-level image features without (or with relatively little) human instruction. These are trained on large image datasets such as Photo.net (datta2006studying), AVA (murray2012ava), and CUHK-PQ (luo2011content), and the specific architecture can vary. Doshi et al. (doshi2020image) fine-tuned two pre-trained models (AlexNet and VGG) to classify images into high and low aesthetic categories. Kao et al. (kao2017deep) investigated using a Multi-Task Convolutional Neutral network (MTCNN) to jointly perform aesthetic assessment and semantic recognition.
@@
In addition to predicting image aesthetics such as quality, beauty, and liking (bartho2023predicting; peng2018feast), there is evidence to suggest that computational aesthetics can serve as a valuable starting point for a variety of downstream tasks.
For instance, one experiment demonstrated the ability to infer the personality traits (“Big Five Inventory”) of Flickr users based on the color, composition, texture, and face features of images that the user tagged as ‘favorites’ on the platform (segalin2016pictures).
Likewise, Machajdik et al. (machajdik2010affective) achieved state-of-the-art results on image emotion classification using a similar set of features, including color, texture, composition, and content, inspired by psychology and art theory. However, results were not uniform: content features such as the number of faces and relative size of the biggest face were good discriminators between fear/disgust vs. amusement in the International Affective Picture System dataset, whereas color features was more predictive in an artistic photography dataset collected from the DeviantArt website.
@@ 2.2 Image Analysis in News Media
One investigation analyzed images of the COVID-19 pandemic appearing in Finnish newspapers, focusing on the social representations of individuals from varying age groups (martikainen2021newspaper). The authors noted that the angle of the photo often contributed to visual rhetoric, for instance, images of children studying at home emphasized a high angle, placing the spectator in the authoritative position overseeing the child whereas those of young adults tended to be taken at eye-level.
Similarly, in a study on how the media covered Pope Francis’ visit to Cuba, Thomson et al. discovered that photo angles in U.S. sources framed the Pope at higher positions than Cuban politicians whereas local sources used a more equal leveling (thomson2018politicians). Another contrast was the content of the images, with Cuban media focusing almost exclusively on the Pope, whereas U.S. media more often published photographs of Cuban people.  (fahmy2010contrasting; rafiee2023framing).
@@
Although these frequently involve analyzing text data to compare sentiments and topics (wallbing2021computational; chen2023leveraging), the visual medium of news media has also been explored. Bhargava et al. (bhargava2020mapping) proposed Photo Spaces, a visualization tool to describe visual narratives for an event, based on image embeddings produced by a pre-trained ResNet-50 model. The spatial positioning of images in the photo-space is determined by reducing these embeddings into two dimensions, via the Uniform Manifold Approximation and Projection technique, from which clusters corresponding to visual themes can be extracted. The authors document an example of the tool (See Figure 1) on abortion rights as a topic, adding the political affinity of sources (right- vs. left-leaning) through the color of each image’s border.
Figure 1: An example of Photo-Space visualization (bhargava2020mapping), where physical placement captures thematic relations, and image borders are color-coded by political affinity.
Similarly, Zhang et al. (zhang2022image) evaluated the effectiveness of bag-of-visual-words, self-supervised learning, and transfer learning to cluster unlabeled image data. They conclude that transfer learning using pre-trained models to obtain embeddings outperforms the other methods, though the choice of the pre-trained model significantly affected clustering results.
@@ 2.3 Video Thumbnails
Thumbnails are small image representations of full-length videos. Thumbnails “tell the users what the video is about (i.e., be informative), and grab potential viewers’ attention (i.e., be visually appealing)” (koh2022exploration).
Kim et al. (kim2023towards) evaluated the design of thumbnails for data-driven news articles, concluding similarly that for thumbnails involving informative graphics, users prefer ones that are interpretable and attractive. For branded content, a thumbnail’s success (measured in view-through) can also be attributed to visual features, such as colorfulness and brightness (koh2022exploration).
@@ 3.1 Data Collection
Inspired by previous work (martikainen2021newspaper), we selected the COVID-19 pandemic as one of the international news events to examine. The complexity of the event gives rise to a diverse set of event themes—medical, economic, political—which may transfer to detectable visual themes. For similar reasons, we also selected the war in Ukraine, a more recent event with different event themes—military conflict, the humanitarian crisis, and geopolitics.
@@ 3.2 Visual Theme Identification
After some experimentation, our best results were obtained by using Concept (grootendorst2024concept), a CLIP and BERTopic-based open source tool that performs topic modeling on images, by first encoding images using CLIP, and then identifying clusters on top of those embeddings. An example of the results produced using this method is shown in Figure 2.
@@
Thumbnail clusters themselves carry no easily interpretable meaning, so to understand the visual themes being extracted we associated each cluster with five textual tags. We leveraged Tag2Text (huang2023tag2text), a model capable of recognizing 3,429 commonly human-used categories, to generate tags for every image. Then, on a cluster level, we aggregated these tags to generate five tags that are most representative of the cluster’s theme.
@@ 3.3 Image Feature Extraction
Image features are extracted using a hand-crafted approach, where various pre-defined statistical measures are computed for each image. An exhaustive reference list of the 21 features, in eight feature groups, is displayed in Table 2. The majority of our feature set is adopted from Bartho et al. (bartho2023predicting). A brief justification of each feature is included here, but the reader is encouraged to consult the references for the computational details and tuning parameters of each algorithm.
@@
Color
(datta2006studying), (li2009aesthetic),
(mallon2014beauty),
(schifanella2015image),
(thieleking2020art),
(iigaya2021aesthetic), (geller2022statistical)
5 Hue mean of hue channel in HSV space
@@
Dimension
(datta2006studying), (iigaya2021aesthetic)
2 Aspect Ratio image width divided by image height (pixels)
@@
Lightness
(schifanella2015image),
(sidhu2018prediction), (mather2018visual),
(iigaya2021aesthetic), (peli1990contrast)
3 Contrast standard deviation of L* channel in L*a*b* space
@@
HOG
(bosch2007representing),
3 Self-Similarity similarity of HOG features
@@
Fourier (redies2008fractal) 2 Fourier Slope
@@
Symmetry
2 Symmetry-LR
@@
CNN (brachmann2017using) 2 Sparseness
@@
Other
(savardi2018shot),
>1 Shot Scale close, medium, or long shot based on fine-tuned CNN
@@
Hue, Saturation, Lab(a), Lab(b), Color Entropy. The use of colors can affect the perceived mood of an image. Although there are multiple ways to represent colors numerically, the HSV space (for hue, saturation, and value), and the L*a*b* space (for lightness, green-red, and blue-yellow) both have been shown to be more aligned with the human perception of color (fadaei2021comparision). Color-entropy captures the visual percept of the “colorfulness” of an image.
@@ 4 Validations
Using the PlacesCNN model (zhou2014learning), which can label each image as indoor/outdoor, we measured the entropy of each theme using the Gini index of statistical dispersion.
@@ 5 Results
Figure 7: Preferences for image aesthetics. Blue: U.S. is greater. Red: Chinese is greater. Bold: statistically significant at $p<.05$.
@@ 7 Application to Disinformation
Previous investigation has shown that multimodal disinformation that uses a text-plus-visual format is perceived as more credible than pure text (hameleers2020picture). Visual disinformation can involve visual content that is deliberately manipulated or generated, or real but used out of context, or real but used in tandem with false claims (cao2020exploring). This paper focuses on the first two.
@@
Ethics Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? Yes
    5. (e)
      Did you describe the limitations of your work? Yes
    6. (f)
      Did you discuss any potential negative societal impacts of your work? No. The work is purely observational.
    7. (g)
      Did you discuss any potential misuse of your work? No. No system was designed or created.
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? No. Data is publicly available from news channels and has not identified individuals without their consent.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? No theoretical results.
    2. (b)
      Have you provided justifications for all theoretical results? No theoretical results.
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? No theoretical results.
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? Yes
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? Yes. Dataset is limited in size.
    6. (f)
      Have you related your theoretical results to the existing literature in social science? No theoretical results.
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? No theoretical results.
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? No theoretical proofs.
    2. (b)
      Did you include complete proofs of all theoretical results? No theoretical proofs.
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? No. No substantive code was created, and the off-the-shelf CNN is cited. Data is publicly available via YouTube API.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? Yes
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? No. Experiments were statistical and run once.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? No. No training of any system was involved.
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? Yes.
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? No. The paper notes preferences and no hard decisions.
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? Yes
    2. (b)
      Did you mention the license of the assets? No. The system (Cluster) and data source (YouTube) are common and well-known.
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? No new assets.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? No consent was required.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? No. Data was standard YouTube download off well-established news channels.
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR (see fair)? No new datasets.
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see gebru2021datasheets)? No new datasets.
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? No human subjects.
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? No human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? No human subjects.
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? No human subjects.